OpenStudy (anonymous):

The line y=-2x+6 meets the x-axis at the point P. the line y=3/2x-4 meets the y-axis at the point Q. find the equation of the line joining the points P and Q

OpenStudy (anonymous):

when y = 0 x = 3 = poin t P on x axis from second equn. y when x = 0, y = -4 = point on y axis Q equation of line PQ: slope = -4 / 3 y-y1 = m(x-x1) y - 0 = -4/3 ( x - 3) y = (-4/3) x + 4